# A Comprehensive Review of Recommender Systems: Transitioning from Theory to Practice

### v2

@@ Abstract
Recommender Systems (RS) play an integral role in enhancing user experiences by providing personalized item suggestions. This survey reviews the progress in RS inclusively from 2017 to 2024, effectively connecting theoretical advances with practical applications. We explore the development from traditional RS techniques like content-based and collaborative filtering to advanced methods involving deep learning, graph-based models, reinforcement learning, and large language models. We also discuss specialized systems such as context-aware, review-based, and fairness-aware RS. The primary goal of this survey is to bridge theory with practice. It addresses challenges across various sectors, including e-commerce, healthcare, and finance, emphasizing the need for scalable, real-time, and trustworthy solutions. Through this survey, we promote stronger partnerships between academic research and industry practices. The insights offered by this survey aim to guide industry professionals in optimizing RS deployment and to inspire future research directions, especially in addressing emerging technological and societal trends111 The ideas and discussions produced in this work are strictly those of the authors and do not represent the points of view of the institutions the authors belong to.. The survey resources are available in the public GitHub repository https://github.com/VectorInstitute/Recommender-Systems-Survey .
Keywords Recommender Systems  $\cdot$ Graph-based Recommender Systems  $\cdot$ Knowledge-based Systems  $\cdot$ Multimodal Recommender Systems  $\cdot$ Large Language Models  $\cdot$ Personalization  $\cdot$ Industry Applications  $\cdot$ Explainable AI  $\cdot$ Transparency  $\cdot$ Fairness  $\cdot$ Deep Learning  $\cdot$ Survey
@@ 1 Introduction
Necessity of this Survey
Previous surveys often concentrate solely on the theoretical aspects of RS, exploring methods and algorithmic foundations to improve prediction accuracy and personalization [13, 4]. Conversely, practical-focused research or applications typically views RS as essential tools for enhancing user engagement, retention, and business growth [14, 15]. There is a need for the collaboration between academia and industry to address both technical challenges and real-world demands, which in turn enhances user satisfaction and business value. This interdependence highlights the growing importance of such partnerships.
Difference with Existing Surveys
Unlike previous surveys that often focus solely on the theoretical or practical aspects, our survey uniquely covers the integration of theoretical advancements with practical applications, offering a comprehensive overview that addresses both academic and industry perspectives. Furthermore, we identify emerging trends and future research directions, such as the integration of explainable AI in RS to ensure transparency and user trust.
@@ 3 Literature Review Methodology
Bibliometric analysis
Figure 1 presents a comprehensive analysis of publications reviewed in this survey. The top left chart shows the distribution of articles by type, highlighting the dominance of experimental studies. The top right chart details the publication trends by year. The bottom left chart lists the top journals, with ’ACM Transactions on Information Systems’ leading in terms of publication count. The bottom middle chart displays the top conferences, with the ’ACM SIGIR Conference on Research and Development in Information Retrieval’ being the most frequented.
@@ 5 Foundational Recommender Systems
A RS can be mathematically represented as a function $f$ that predicts the utility of an item $i$ for a user $u$, denoted as $\hat{r}_{ui}$, which estimates how much user $u$ would prefer item $i$.[3] . This function is typically learned from historical data:
@@ 5.1 Foundational Recommender Systems
Algorithmic advancements in computer science and ML methods led to a shift to sophisticated neural networks that allow automatic feature extraction and learning complex data patterns [86]. These neural network-based systems leverage deep learning to analyze user interactions and item features across different modalities, including textual (such as reviews [87], citations [88], and news [89]), streaming (like music [77]), and image data [90]. The overall goal of these advancements is to generate personalized recommendations by aligning user profiles with item characteristics.
@@
  • $I$ is a $k\times n$ matrix, with $n$ being the number of items.
Neural extensions in CF have advanced these RS, utilizing deep learning to capture intricate user-item relationships and significantly improve recommendation accuracy. Techniques such as Neural Collaborative Filtering (NCF) [94], Sequence-Aware RS [95], and Graph Neural Networks (GNNs) [48] have emerged as state-of-the-art approaches in CF.
@@ 6.1 Deep Learning-based Recommender Systems
Multi-Layer Perceptrons Traditional RS primarily use linear methods like matrix factorization [92], which struggle with capturing complex user-item interactions. In contrast, Multi-Layer Perceptrons (MLPs), a type of feedforward neural network, use deep layers to model these nonlinear interactions more accurately, improving both prediction accuracy and recommendation quality. The evolution of MLPs is seen in RS such as Neural Collaborative Filtering (NCF) [94], Deep Factorization Machine (DFM) [99], Wide & Deep [97], xDeepFM [141], Deep & Cross Network (DCN) [100], FMLP-Rec [142], a model with learnable filters for improving sequential recommendation—and FinalMLP [143], which combines dual MLP architectures with feature selection for effective Click-Through Rate (CTR) prediction.
@@
Notable RS include AutoRec [144], Collaborative Filtering Autoencoder, Multi-Variational Auto-Encoder (Multi-VAE) [145], Deep Recommender (DeepRec) [146], Recommender Variational Auto-Encoder (RecVAE) [147], Item-based variational auto-encoder for fair recommendation [148], and the Variational Bandwidth Auto-Encoder (VBAE) hybrid RS [149] . These approaches address sparsity and noise challenges, making them effective for personalized recommendations.
@@
Convolutional Neural Networks (CNNs) can learn from visual, sequential, and multimodal data and have enhanced accuracy and personalization of recommendations. CNNs have been applied in RS in various settings. DeepCoNN analyzes text and visual cues to understand user preferences [87, 151]. CNNs are integrated with graph structures for scalable recommendation systems [120], employed in DKN for news recommendations [152], and utilized in MusicCNN for music recommendations based on audio signals [153]. CNN-based RS models predict next-item recommendations [130], recognize user preference patterns through CoCNN [154], and leverage collaborative filtering with CAGCN [155].
@@
Recurrent Neural Networks (RNNs) are adept at capturing complex user-item interactions within sequential data [157]. The evolution of RNN-based RS began with GRU4Rec, utilizing Gated Recurrent Units (GRUs) for session-based recommendations [158, 159]. NARM introduced an attention mechanism to enhance accuracy [160], while SASRec used self-attention to capture long-term semantics [161]. Deep attention neural networks were employed for session-based recommendations [162]. RNNs, including LSTMs and GRUs, comprehend temporal dynamics in user behavior [163, 164]. Integrating RNNs with CNNs via recurrent convolutional networks offers deeper insights into user preferences [165], followed by CNN-RNN hybrid RS [166] and Reinforcement Learning-based cross-domain recommendations [167]. Additionally, a knowledge graph recommendation algorithm using RNN encoders has emerged [168].
@@ 7.1 Graph-based Recommender Systems
The introduction of LightGCN [202], with its focus on neighborhood aggregation and streamlined architecture, represents a simplification in the GNN landscape, improving efficiency without compromising performance. MixGCF [203] brought forward a novel approach to negative sampling, optimizing training processes. Subsequent developments like GNNRec [187] and XSimGCL [222] advanced session-based and graph contrastive learning recommendations, respectively, addressing specific challenges such as social influence integration and bias mitigation. Ensuring trustworthiness in GNN-based RS requires enhancements in robustness, explainability, and fairness to ensure reliable recommendations [224].
@@ 7.2 Sequential and Session-based Recommender Systems
Frequency Enhanced Hybrid Attention Network (FEARec) [240] and Knowledge Prompt-tuning for Sequential Recommendation (KP4SR) [241] advance sequential recommendation by leveraging hybrid attention mechanisms and integrating external knowledge bases, respectively, for better model performance.
@@ 7.3 Knowledge-based Recommender Systems
Propagation-based approaches influence embeddings through multi-hop neighbor interactions within the KG. Item KG-based methods like Ripplenet [215] aggregate item-related embeddings to derive user interests, whereas User-Item KG-based methods such as KGAT [197] and Intentgc [195] refine both user and item embeddings by propagating embeddings across a user-item graph, enhancing recommendation accuracy.
@@ 7.4 Reinforcement Learning-based Recommender Systems
Furthermore, model-based RL approaches in RS focus on simulating user behavior to tailor recommendations, with techniques ranging from generative adversarial networks [119] to multi-agent systems [260, 268]. These sophisticated methods aim to predict user interactions and refine recommendations continually, enhancing personalization and contextual relevance in RS.
@@
Table 9: Comprehensive Overview of Reinforcement Learning based Recommender System Models across Various Metrics and Use Cases. This table details each model’s Input features, Year of Publication, and Characteristics such as Scalability, Interpretability, Efficiency, and Reproducibility (rated as High, Medium, or Low). It also lists the Dataset Used, Evaluation Metrics, Model Accuracy, Learning Task, and Application Field. Metrics that their numerical value is not reported are specified with “No numerical value”.
@@
LIRD[252] 2017 User-item graphs; node features -, -, -, Low E-commerce website MAP, NDGC No numerical value e-commerce
@@
[271] 2018 User-item graphs; node features -, Low, Medium, Low Movielens-100k, Movielens-1M P@30, R@30 Movielens-100k: 0.246/0.169 Movielens-1M: 0.277/0.155 e-commerce
[119] 2018 User-item graphs; node features -, Medium, -, High MovieLens, LastFM, Yelp, Taobao, YooChoose, Ant Financial Reward, CTR Combined datset: 25.36/0.78161616Shown only the best results e-commerce
@@
KERL[230] 2020 User-item graphs; node features, KG -, -, -, Low Beauty, CD, Books, LastFM HR@10, NDCG@10 Beauty: 54.1/36.5 CD: 73.7/50.8 Books: 80.0/57.1 LastFM: 64.2/50.1  [^17] Shown next-item recommendation metrics e-commerce
@@
TRIGR[217] 2022 User-item graphs; node & item features Medium, -, Medium, High Music, Beauty, Clothing HR@10, F1@10, NDGC@10 Music: 0.9886/0.2304/ 0.9436 Beauty: 0.8845/0.1798/ 0.6949 Clothing: 0.7544/0.1405/ 0.4865 e-commerce
@@
MDP[257] 2021 User-item interactions -, -, -, Low Transactional data of job applications % improvement of wage, market revenue, worker success measures 22%, 1.5-6%, 4x e-learning
@@ 7.5 Large Language Model based Recommender Systems
Moreover, advancements in prompt tuning and personalized recommendation strategies demonstrate the ongoing evolution of LLM applications in RS, significantly improving system performance while also highlighting challenges such as ethical considerations and the management of popularity biases [286, 283]. These developments indicate a move towards more sophisticated, context-aware systems that can dynamically adapt to user preferences and behaviors.
@@
MLPR [347] 2022 User-item graphs; node features High,Medium, High, High One month data from Walmart.com AUC; NDCG@1242424NDCG@5 was omitted. Click: +6.48%, +17.22%; ATC: +4.66%, +10.61%; Purchase: +1.03%, +5.36% E-commerce
@@
P5 [345] 2022 User-item graphs; node features High, Medium,High,High [^28] https://github.com/jeykigung/P5 Amazon (Sports, Beauty, Toys), Yelp For Performance on Sequential Recommendations [^29] We dropped other details about performance on rating, explanation generation and review preference, and considered only the performance comparison on sequential recommendation because its the most relevant factor in this case.: HR@10, NDCG@10303030HR@5 and NDCG@5 are dropped for simplicity.; For P5-B [^31] Only P5-base scenario is considered for simplicity. Amazon Sports: 0.0460, 0.0336; Amazon Beauty: 0.0.0645, 0.0416; Amazon Toys: 0.0675, 0.0536 E-commerce
RecMind [293] 2024 User-item graphs; node features High, Medium, High,Low Amazon Reviews - Beauty; Yelp For Performance on Sequential Recommendations [^32] We dropped other details about performance on rating, explanation generation and review preference, and considered only the performance comparison on sequential recommendation because its the most relevant factor in this case.: HR@10, NDCG@10333333HR@5 and NDCG@5 are dropped for simplicity. For RecMind-SI (few-shot) [^34] Only RecMind-SI (few-shot) scenario is considered for simplicity and its high performance. Amazon Reviews - Beauty: 0.1559, 0.1063; Yelp: 0.2451, 0.1607 E-commerce; Restaurants

### v4

@@ Abstract
Recommender Systems (RS) play an integral role in enhancing user experiences by providing personalized item suggestions. This survey reviews the progress in RS inclusively from 2017 to 2024, effectively connecting theoretical advances with practical applications. We explore the development from traditional RS techniques like content-based and collaborative filtering to advanced methods involving deep learning, graph-based models, reinforcement learning, and large language models. We also discuss specialized systems such as context-aware, review-based, and fairness-aware RS. The primary goal of this survey is to bridge theory with practice. It addresses challenges across various sectors, including e-commerce, healthcare, and finance, emphasizing the need for scalable, real-time, and trustworthy solutions. Through this survey, we promote stronger partnerships between academic research and industry practices. The insights offered by this survey aim to guide industry professionals in optimizing RS deployment and to inspire future research directions, especially in addressing emerging technological and societal trends. The survey resources are available in the public GitHub repository https://github.com/VectorInstitute/Recommender-Systems-Survey .
keywords:
Recommender Systems , Graph-based Recommender Systems , Knowledge-based Systems , Multimodal Recommender Systems , Large Language Models , Personalization , Industry Applications , Explainable AI , Transparency , Fairness , Deep Learning , Survey
[…]
[inst1]organization=Vector Institute, addressline=, city=Toronto, state=ON, postcode=, country=Canada
[…]
[inst2]organization=Independent Researcher, addressline=, city=Toronto, state=ON, postcode=, country=Canada
DisclaimerDisclaimerfootnotetext: The ideas and discussions produced in this work are strictly those of the authors and do not represent the points of view of the institutions the authors belong to.
@@ 1 Introduction
Necessity of this Survey Previous surveys often concentrate solely on the theoretical aspects of RS, exploring methods and algorithmic foundations to improve prediction accuracy and personalization [13, 4]. Conversely, practical-focused research or applications typically views RS as essential tools for enhancing user engagement, retention, and business growth [14, 15]. There is a need for the collaboration between academia and industry to address both technical challenges and real-world demands, which in turn enhances user satisfaction and business value. This interdependence highlights the growing importance of such partnerships.
Difference with Existing Surveys Unlike previous surveys that often focus solely on the theoretical or practical aspects, our survey uniquely covers the integration of theoretical advancements with practical applications, offering a comprehensive overview that addresses both academic and industry perspectives. Furthermore, we identify emerging trends and future research directions, such as the integration of explainable AI in RS to ensure transparency and user trust.
@@ 3 Literature Review Methodology
Figure 1 presents a comprehensive analysis of publications reviewed in this survey.
@@ 5 Foundational Recommender Systems
A RS can be mathematically represented as a function $f$ that predicts the utility of an item $i$ for a user $u$, denoted as $\hat{r}_{ui}$, which estimates how much user $u$ would prefer item ii.[3] . This function is typically learned from historical data:
@@ 5.1 Foundational Recommender Systems
Algorithmic advancements in computer science and ML methods led to a shift to sophisticated neural networks that allow automatic feature extraction and learning complex data patterns [86]. These neural network-based systems leverage deep learning to analyze user interactions and item features across different modalities, including textual (such as reviews [87], citations [88], and news [89]), streaming (like music [77]), and image data [90]. In practice, the tunable choices are the feature-weighting/normalization scheme and the similarity metric (e.g., cosine or Pearson), whereas any similarity cutoff for top-N selection is an operational post-hoc threshold rather than a model hyperparameter [91, 92]. The overall goal of these advancements is to generate personalized recommendations by aligning user profiles with item characteristics.
@@
  • 2.
    II is a $k\times n$ matrix, with $n$ being the number of items.
Neural extensions in CF have advanced these RS, utilizing deep learning to capture intricate user-item relationships and significantly improve recommendation accuracy. Techniques such as Neural Collaborative Filtering (NCF) [96], Sequence-Aware RS [97], and Graph Neural Networks (GNNs) [48] have emerged as state-of-the-art approaches in CF. For MF-based CF we tune latent dimension, L2 regularization, and learning rate to balance expressiveness and overfitting [98], while in neighborhood-based CF the neighborhood size is an operational setting and should be discussed with memory-based methods [99].
@@ 6.1 Deep Learning-based Recommender Systems
Multi-Layer Perceptrons Traditional RS primarily use linear methods like matrix factorization [94], which struggle with capturing complex user-item interactions. In contrast, Multi-Layer Perceptrons (MLPs), a type of feedforward neural network, use deep layers to model these nonlinear interactions more accurately, improving both prediction accuracy and recommendation quality. The evolution of MLPs is seen in RS such as Neural Collaborative Filtering (NCF) [96], Deep Factorization Machine (DFM) [103], Wide & Deep [101], xDeepFM [145], Deep & Cross Network (DCN) [104], FMLP-Rec [146], a model with learnable filters for improving sequential recommendation—and FinalMLP [147], which combines dual MLP architectures with feature selection for effective Click-Through Rate (CTR) prediction. MLP recommenders typically select embedding size, hidden-layer count, and dropout via validation to trade off capacity and generalization [148, 103].
@@
Notable RS include AutoRec [149], Collaborative Filtering Autoencoder, Multi-Variational Auto-Encoder (Multi-VAE) [150], Deep Recommender (DeepRec) [151], Recommender Variational Auto-Encoder (RecVAE) [152], Item-based variational auto-encoder for fair recommendation [153], and the Variational Bandwidth Auto-Encoder (VBAE) hybrid RS [154] . These approaches address sparsity and noise challenges, making them effective for personalized recommendations. Vanilla AE recommenders tune the latent (bottleneck) size, learning rate, and L2 weight decay (or dropout), with reconstruction as the primary objective; only multi-objective variants introduce method-specific loss weights [149, 155].
@@
Convolutional Neural Networks (CNNs) can learn from visual, sequential, and multimodal data and have enhanced accuracy and personalization of recommendations. CNNs have been applied in RS in various settings. DeepCoNN analyzes text and visual cues to understand user preferences [87, 157]. CNNs are integrated with graph structures for scalable recommendation systems [124], employed in DKN for news recommendations [158], and utilized in MusicCNN for music recommendations based on audio signals [159]. CNN-based RS models predict next-item recommendations [134], recognize user preference patterns through CoCNN [160], and leverage collaborative filtering with CAGCN [161].For CNN-based recommenders, kernel size, stride, and filter count control the granularity of local pattern capture (distinct from two-tower MLP systems such as YouTube’s) [162, 163, 164]. For CNN-based recommenders, kernel size, stride, and filter count control the granularity of local pattern capture [162, 163, 164].
@@
Recurrent Neural Networks (RNNs) are adept at capturing complex user-item interactions within sequential data [166]. The evolution of RNN-based RS began with GRU4Rec, utilizing Gated Recurrent Units (GRUs) for session-based recommendations [167, 168]. NARM introduced an attention mechanism to enhance accuracy [169], while SASRec used self-attention to capture long-term semantics [170]. Deep attention neural networks were employed for session-based recommendations [171]. RNNs, including LSTMs and GRUs, comprehend temporal dynamics in user behavior [172, 173]. Integrating RNNs with CNNs via recurrent convolutional networks offers deeper insights into user preferences [174], followed by CNN-RNN hybrid RS [175] and Reinforcement Learning-based cross-domain recommendations [176]. Additionally, a knowledge graph recommendation algorithm using RNN encoders has emerged [177]. RNN sequential recommenders (e.g., GRU4Rec) typically tune hidden-state size, maximum sequence length, and dropout to balance temporal coverage and training stability [178, 179].
@@
Self-supervised learning (SSL)–based recommenders. SSL introduces auxiliary objectives that derive supervision from unlabeled user–item interactions [180]. Contrastive-learning–based models include SGL [181] graph view augmentations via node and edge dropout for robustness under sparsity, SimGCL [182] noise-perturbed embedding views without explicit graph ops, LightGCL [183] lightweight propagation with SVD-guided signal, NCL [184] neighborhood-enriched positives, HCCF [185] hypergraph views for higher-order relations, KGCL [186] adapts SSL to knowledge graphs via cross-view alignment between CF and KG signals, AdaGCL [187] learns augmentation strengths with trainable generators, CL4SRec [188] masking/cropping/segment permutation for sequence views, TGCL [189] time-aware transitions with temporal contrast, and IDCL [190] intent-disentangled graph contrastive learning. Generative/predictive SSL includes BERT4Rec [191] masked-item prediction, S3-Rec [192] multi-task mutual-information objectives over items, attributes, and subsequences, and MAERec [193] masked autoencoding for recommendation. Other/hybrid SSL directions feature BUIR [194] BYOL-style bootstrap without negatives, DuoRec [195] dual contrastive objectives to improve sequence representations, and MMGCL [196] multimodal contrastive learning integrating text and vision with interactions.
Key SSL hyperparameters include the temperature $\tau$ in the InfoNCE family (controls similarity sharpness), the number and sampling strategy of negatives (influences hardness and informativeness of comparisons), and augmentation probabilities (e.g., edge-drop or masking ratios) that govern view diversity. Their tuning primarily shapes the alignment–uniformity trade-off in learned embeddings [197, 180].
Challenges with SSL: SSL-based recommenders remain sensitive to augmentation design, negative sampling, and temperature tuning; weak or excessive perturbations can cause representation collapse or semantic drift, and optimal settings are often dataset-specific [180, 197, 195].
@@ 7.1 Graph-based Recommender Systems
The introduction of LightGCN [231], with its focus on neighborhood aggregation and streamlined architecture, represents a simplification in the GNN landscape, improving efficiency without compromising performance. MixGCF [232] brought forward a novel approach to negative sampling, optimizing training processes. Subsequent developments like GNNRec [216] and XSimGCL [251] advanced session-based and graph contrastive learning recommendations, respectively, addressing specific challenges such as social influence integration and bias mitigation. Ensuring trustworthiness in GNN-based RS requires enhancements in robustness, explainability, and fairness to ensure reliable recommendations [253]. Graph-based recommenders adjust the number of propagation layers, neighborhood-sampling size, and embedding dimension to capture higher-order connectivity without oversmoothing [334, 231].
@@ 7.2 Sequential and Session-based Recommender Systems
Frequency Enhanced Hybrid Attention Network (FEARec) [268] and Knowledge Prompt-tuning for Sequential Recommendation (KP4SR) [269] advance sequential recommendation by leveraging hybrid attention mechanisms and integrating external knowledge bases, respectively, for better model performance. Transformer-style sequential models (e.g., SASRec, BERT4Rec) tune hidden size, number of self-attention heads, number of layers (stack depth), and maximum sequence length to set the temporal range and computational cost [338, 191].
@@ 7.3 Knowledge-based Recommender Systems
Propagation-based approaches influence embeddings through multi-hop neighbor interactions within the KG. Item KG-based methods like Ripplenet [244] aggregate item-related embeddings to derive user interests, whereas User-Item KG-based methods such as KGAT [226] and Intentgc [224] refine both user and item embeddings by propagating embeddings across a user-item graph, enhancing recommendation accuracy. For knowledge-graph–based recommenders (as opposed to symbolic rule systems), key hyperparameters include embedding dimension, number of message-passing layers, and regularization strength [226, 344].
@@ 7.4 Reinforcement Learning-based Recommender Systems
Furthermore, model-based RL approaches in RS focus on simulating user behavior to tailor recommendations, with techniques ranging from generative adversarial networks [123] to multi-agent systems [288, 296]. These sophisticated methods aim to predict user interactions and refine recommendations continually, enhancing personalization and contextual relevance in RS. RL-based recommenders commonly tune the discount factor $\gamma$ (short- vs. long-term reward trade-off), exploration rate $\varepsilon$, learning rate, and target-update schedule for stability [360, 361, 362].
@@
Table 9: Comprehensive Overview of Reinforcement Learning based Recommender System Models across Various Metrics and Use Cases. This table details each model’s Input features, Year of Publication, and Characteristics such as Scalability, Interpretability, Efficiency, and Reproducibility (rated as High, Medium, or Low). It also lists the Dataset Used, Evaluation Metrics, Model Accuracy, Learning Task, and Application Field. Metrics that do not report numerical values are marked “No numerical value”.
@@
LIRD [280] 2017 User-item graphs; node features -, -, -, Low E-commerce website MAP, NDCG No numerical value e-commerce
@@
[299] 2018 User-item graphs; node features -, Low, Medium, Low MovieLens-100k, MovieLens-1M P@30, R@30 ML-100k: 0.246/0.169 ML-1M: 0.277/0.155 e-commerce
[123] 2018 User-item graphs; node features -, Medium, -, High MovieLens, LastFM, Yelp, Taobao, YooChoose, Ant Financial Reward, CTR Combined dataset: 25.36/0.78181818Shown only the best results e-commerce
@@
KERL [258] 2020 User-item graphs; node features, KG -, -, -, Low Beauty, CD, Books, LastFM HR@10, NDCG@10 Beauty: 54.1/36.5 CD: 73.7/50.8 Books: 80.0/57.1 LastFM: 64.2/50.1191919Shown next-item recommendation metrics e-commerce
@@
TRIGR [246] 2022 User-item graphs; node & item features Medium, -, Medium, High Music, Beauty, Clothing HR@10, F1@10, NDCG@10 Music: 0.9886/0.2304/ 0.9436 Beauty: 0.8845/0.1798/ 0.6949 Clothing: 0.7544/0.1405/ 0.4865 e-commerce
@@
MDP [285] 2021 User-item interactions -, -, -, Low Transactional data of job applications % improvement of wage, market revenue, worker success measures 22%, 1.5–6%, 4x e-learning
@@ 7.5 Large Language Model based Recommender Systems
Moreover, advancements in prompt tuning and personalized recommendation strategies demonstrate the ongoing evolution of LLM applications in RS, significantly improving system performance while also highlighting challenges such as ethical considerations and the management of popularity biases [314, 311]. These developments indicate a move towards more sophisticated, context-aware systems that can dynamically adapt to user preferences and behaviors. LLM recommenders adjust generation hyperparameters (temperature, top-pp/top-$k$) and fine-tuning optimizers (e.g., learning rate/weight decay), with PEFT methods introducing method-specific knobs such as LoRA rank; prompt length is constrained by the context window [381, 382, 383, 384].
@@
MLPR [385] 2022 User-item graphs; node features High,Medium, High, High One month data from Walmart.com AUC; NDCG@1262626NDCG@5 was omitted. Click: +6.48%, +17.22%; ATC: +4.66%, +10.61%; Purchase: +1.03%, +5.36% E-commerce
@@
P5 [379] 2022 User-item graphs; node features High, Medium,High,High [^30] https://github.com/jeykigung/P5 Amazon (Sports, Beauty, Toys), Yelp For Performance on Sequential Recommendations [^31] We dropped other details about performance on rating, explanation generation and review preference, and considered only the performance comparison on sequential recommendation because its the most relevant factor in this case.: HR@10, NDCG@10323232HR@5 and NDCG@5 are dropped for simplicity.; For P5-B [^33] Only P5-base scenario is considered for simplicity. Amazon Sports: 0.0460, 0.0336; Amazon Beauty: 0.0.0645, 0.0416; Amazon Toys: 0.0675, 0.0536 E-commerce
RecMind [321] 2024 User-item graphs; node features High, Medium, High,Low Amazon Reviews - Beauty; Yelp For Performance on Sequential Recommendations [^34] We dropped other details about performance on rating, explanation generation and review preference, and considered only the performance comparison on sequential recommendation because its the most relevant factor in this case.: HR@10, NDCG@10353535HR@5 and NDCG@5 are dropped for simplicity. For RecMind-SI (few-shot) [^36] Only RecMind-SI (few-shot) scenario is considered for simplicity and its high performance. Amazon Reviews - Beauty: 0.1559, 0.1063; Yelp: 0.2451, 0.1607 E-commerce; Restaurants